### v1

# The exact travelling wave solutions of a KPP equation

Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. The solutions are written in terms of elementary functions.
I Introduction
Fisher fisher first introduced a nonlinear evolution equation to investigate the wave propagation of an advantageous gene in a population. His equation also describes the logistic growth–diffusion process and has the form
$u_{t}-\nu u_{xx}=ku\left(1-\frac{u}{\kappa}\right)$ (1)
where $\kappa(>0)$ is the carrying capacity of the environment. The term $f(u)=ku(1-u/\kappa)$ represents a nonlinear growth rate which is proportional to $u$ for small $u$, but decreases as $u$ increases, and vanishes when $u=\kappa$. Equation (1) arises in many physical, biological, and chemical problems involving diffusion and nonlinear growth. It is convenient to introduce the nondimensional quantities $x^{*},t^{*},u^{*}$, so after dropping the asterisks, equation (1) takes the nondimensional form
$u_{t}-u_{xx}=u\left(1-u\right).$ (2)
Fisher equation admits traveling wave solutions
$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (3)
where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=u\left(u-1\right).$ (4)
The aim of the present paper is to show that the generalization of Eq. (4), to be presented below, can be integrated in elementary functions for the specific values of the speed $c$.
In the same year (1937) as Fisher, Kolmogorov, Petrovsky and Piskunov (KPP) kolmogorov introduced the more general reaction-diffusion equation, which in the nondimensional form can be presented as
$u_{t}-u_{xx}=F(u),$ (5)
where $F$ is a sufficiently smooth function with the properties that $F(0)=F(1)=0$, $F^{\prime}(0)>0$ and $F^{\prime}(1)<0$.
II Travelling waves
We consider the equation
$u_{t}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (6)
where $n$ and $a$ are some positive constants. Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (6) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath . For the solutions of (6) in the form (3), $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=u\left(u^{n}-1\right)\left(u^{n}+a\right)$ (7)
with the boundary conditions (11). Equation (7) has a simple mechanical interpretation. It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
$\frac{dU}{du}=-u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (8)
The motion is with friction, and the coefficient of friction is equal to $c$. The point $u=0$ is the point of stable equilibrium, and the point $u=1$ unstable.
Equation (7) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
$p=\frac{du}{d\xi}$ (9)
as the new dependent variable. In the new variables (7) takes the form of Abel equation of the second kind polyanin .
$pp_{u}+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (10)
The boundary conditions for (7) were
$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (11a)
$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (11b)
The boundary conditions for (10) are
$p(1)$ $=1,$ (12a)
$p(0)$ $=0.$ (12b)
One can easily check up that for
$c=\frac{(n+1)a+1}{\sqrt{n+1}},$ (13)
the solution of (10) satisfying the boundary conditions (12) is
$p=\frac{u\left(u^{n}-1\right)}{\sqrt{n+1}}.$ (14)
Substituting $p(u)$ into (9) and integrating we obtain the solution of (7) as
$u(\xi)=\frac{1}{\left[\exp\left(n\xi/\sqrt{n+1}\right)+1\right]^{1/n}}.$ (15)
For the Fisher equation we obtain the well known result ablo
$u(\xi)=\frac{1}{\left[\exp\left(\xi/\sqrt{6}\right)+1\right]^{2}},$ (16)
and the condition (13) takes the form $c=5/\sqrt{6}$.
III Perturbation series
In the previous Section we used the method of integration based on Eq. (10) to find the exact integral of Eq. (7) for the specific value of $c$. Now we want to consider the limiting case $c\gg 1$ and show how the former equation can be used to construct the perturbation series for the solution of the latter.
Expanding the solution of (10) with respect to the parameter $1/c$
$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (17)
we obtain
$p^{(1)}$ $=u\left(u^{n}-1\right)\left(u^{n}+a\right),$ (18a)
$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (18b)
To make our life simpler, further on consider the particular case of (10) corresponding to the Fisher equation:
$pp_{u}+cp=u(u-1).$ (19)
In this case the first two terms of the expansion (17) are
$p^{(1)}$ $=u(u-1),$ (20a)
$p^{(2)}$ $=u(u-1)(1-2u).$ (20b)
Substituting $p=p^{(1)}+p^{(2)}/c$ into (9) we obtain after a bit of algebra debnath
$u(\xi)$ $=$ $\frac{1}{\exp\left(\xi/c\right)+1}$ (21)
$+$ $\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$
IV Discussion
Let us return to (10). We have the first order ODE with two boundary conditions (12). The first impression is that the problem is overdetermined, and the solution obtained in Section II exists due to the finetuning of the l.h.s. of the equation (by choosing $c$ given by Eq. (13)). (It is irrelevant for the time being that the solution both of (10) and (7) turn out to be the elementary functions.) And the solution of (10) will not exist for any other value of $c$. However, the last statement immediately comes into contradiction with the results of the previous Section. So what about our first impression?
Notice that the boundary conditions (12) for Eq. (10) are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$. So the boundary conditions should be taken with care. In what follows, we’ll understand that all boundary conditions are equal, but some are more equal than the others.
Let us return to Eq. (7) and consider the solution in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
$u_{\xi}$ $=p,$ (22a)
$p_{\xi}$ $=-cp-au.$ (22b)
In the second case, the equation can be presented as
$v_{\xi}$ $=-p,$ (23a)
$p_{\xi}$ $=cp-(a+1)nv,$ (23b)
where $v=1-u$.
Let us start from analysing Eq. (23). The characteristic equation for the system is
$s^{2}-cs-n(a+1)=0,$ (24)
the roots being
$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4n(a+1)}}{2}.$ (25)
(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.) Hence the system can be reduced to the canonical form
$X_{\xi}$ $=s_{1}X,$ (26a)
$Y_{\xi}$ $=s_{2}Y,$ (26b)
where
$X$ $=p+s_{2}v,$ (27a)
$Y$ $=p+s_{1}v.$ (27b)
Because $u(\xi)$ should satisfy the boundary condition (11a), we obtain
$Y=0.$ (28)
Returning to Eq. (10) we realize that the boundary condition (12a) means that in the vicinity of $u=1$
$p(u)=s_{1}(u-1).$ (29)
Now let us come to (22). The characteristic equation for the system is
$s^{2}+cs+a=0.$ (30)
The roots of the characteristic equations are
$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4a}}{2}.$ (31)
We have to consider separately the case of the real roots, the case of the complex roots and the case of the double root.
Let us start from the case of the real roots ($c>2\sqrt{a}$). In this case we can reduce the system (22) to
$\frac{dX}{s_{3}X}=\frac{dY}{s_{4}Y},$ (32)
where
$X$ $=p+s_{4}v,$ (33a)
$Y$ $=p+s_{3}v.$ (33b)
The general integral is given by the equation
$Y=CX^{\frac{s_{2}}{s_{1}}}.$ (34)
Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node. (Note that $c$ given by (13) obviously fulfills the inequality $c>2\sqrt{a}$ for any $n>0$ and $a>0$.)
For $c<2\sqrt{a}$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (22) to the form goursat
$\frac{dX}{\alpha X-\beta Y}=\frac{dY}{\beta X+\alpha Y}$ (35)
from which we derive
$\frac{XdX+YdY}{\alpha\left(X^{2}+Y^{2}\right)}=\frac{XdY-YdX}{\beta\left(X^{2}+Y^{2}\right)}$ (36)
The general integral of the equation (122) is therefore represented by the equation
$\sqrt{X^{2}+Y^{2}}=Ce^{\frac{\alpha}{\beta}\tan^{-1}\frac{Y}{X}}.$ (37)
All these curves have the form of spirals which approach the origin as an asymptotic point. The origin is said to he a focus.
If the characteristic equation has a double root $s=c/2$, the system (22) reduces to the form
$\frac{dX}{X}=\frac{dY}{X+Y},$ (38)
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus.
Now we can resolve the (seeming) contradiction mentioned in the beginning of the present Section. Actually we have to solve ODE with a single boundary condition (12a) (presented as (29)). The second boundary condition (12b) is satisfied automatically