### v2

# The exact travelling wave solutions of a KPP equation

Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. The solutions are written in terms of elementary functions.
I Introduction
Fisher fisher first introduced a nonlinear evolution equation to investigate the wave propagation of an advantageous gene in a population. His equation also describes the logistic growth–diffusion process and has the form
$u_{t}-\nu u_{xx}=ku\left(1-\frac{u}{\kappa}\right)$ (1)
where $\kappa(>0)$ is the carrying capacity of the environment. The term $f(u)=ku(1-u/\kappa)$ represents a nonlinear growth rate which is proportional to $u$ for small $u$, but decreases as $u$ increases, and vanishes when $u=\kappa$. Equation (1) arises in many physical, biological, and chemical problems involving diffusion and nonlinear growth. It is convenient to introduce the nondimensional quantities $x^{*},t^{*},u^{*}$, so after dropping the asterisks, equation (1) takes the nondimensional form
$u_{t}-u_{xx}=u\left(1-u\right).$ (2)
In the same year (1937) as Fisher, Kolmogorov, Petrovsky and Piskunov (KPP) kolmogorov introduced the more general reaction-diffusion equation, which in the nondimensional form can be presented as
$u_{t}-u_{xx}=F(u),$ (3)
where $F$ is a sufficiently smooth function with the properties that $F(0)=F(1)=0$, $F^{\prime}(0)>0$ and $F^{\prime}(1)<0$.
II Travelling waves
Equation (3) admits traveling wave solutions
$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (4)
where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=F(u)$ (5)
with the boundary conditions
$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (6a)
$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (6b)
Equation (5) has a simple mechanical interpretation. It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
$\frac{dU}{du}=-F(u).$ (7)
The motion is with friction, and the coefficient of friction is equal to $c$. The point $u=0$ is the point of stable equilibrium, and the point $u=1$ unstable.
III The ODE which doesn’t contain explicitly the independent variable
Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea kogan3 ; kogan4 ; kogan5 to consider $u$ as the new independent variable and
$p=\frac{du}{d\xi}$ (8)
as the new dependent variable. In the new variables (5) takes the form of Abel equation of the second kind polyanin .
$pp_{u}+cp=F(u).$ (9)
The boundary conditions for (5) were (6). The boundary conditions for (9) are (however, see the next Section)
$p(1)$ $=1,$ (10a)
$p(0)$ $=0.$ (10b)
Now let us be more specific and consider $F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right)$, where $n>0$ and $a\geq 0$ are some constants, so Eqs. (5) and (9) respectively take the form
$u_{\xi\xi}+cu_{\xi}$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (11a)
$pp_{u}+cp$ $=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (11b)
Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (11a) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath . One can easily check up that there exists the solution of (11b) satisfying the boundary conditions (10)
$p=Pu\left(u^{n}-1\right),$ (12)
where
$P=\frac{1}{\sqrt{n+1}},$ (13)
provided that the speed of the wave is
$c=P+\frac{a}{P}.$ (14)
Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (11a) as
$u(\xi)=\frac{1}{\left[\exp(nP\xi)+1\right]^{1/n}}.$ (15)
For the Fisher equation we obtain the well known result ablo
$u(\xi)=\frac{1}{\left[\exp\left(\xi/\sqrt{6}\right)+1\right]^{2}},$ (16)
and the condition (14) takes the form $c=5/\sqrt{6}$.
In Appendix A we’ll see how the results (12) - (15) can be generalized in the case when advection is taken into account in addition to diffusion.
IV The correct boundary conditions for Eq. (9)
In the previous Section we postulated two boundary conditions (10) for (9). However, the equation is of the first order. The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s. of the equation. So one may think that we should finetune the speed $c$ for the solution of (11b) to exist, and the solution will not exist for a general value of $c$. In fact, didn’t we finetune $c$ in the previous Section?
However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (11b) for arbitrary $c\gg 1$ is presented. It looks like we have a problem. We have another problem with the boundary conditions (10) - they are exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$. In this Section we’ll solve both problems. We’ll show that one of the boundary conditions (10) should be specified, and the other can be discarded.
Consider the solution of Eq. (5) in the vicinity of $u=0$ and $u=1$. In the first case, the equation can be presented as
$u_{\xi}$ $=p,$ (17a)
$p_{\xi}$ $=-cp-F^{\prime}(0)u.$ (17b)
In the second case, the equation can be presented as
$v_{\xi}$ $=-p,$ (18a)
$p_{\xi}$ $=cp-|F^{\prime}(1)|v,$ (18b)
where $v=1-u$.
Let us start from analysing (18). The system can be reduced to the canonical form
$X_{\xi}$ $=s_{1}X,$ (19a)
$Y_{\xi}$ $=s_{2}Y,$ (19b)
where
$X$ $=p+s_{2}v,$ (20a)
$Y$ $=p+s_{1}v,$ (20b)
and $s_{1,2}$ are the roots of the characteristic equation
$s^{2}-cs-|F^{\prime}(1)|=0,$ (21)
that is
$s_{1,2}$ $=\frac{c\pm\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (22)
(Further on $s_{1}$ will be the positive root, and $s_{2}$ the negative.)
Because $u(\xi)$ should satisfy the boundary condition (6a), we obtain
$Y=0.$ (23)
Returning to Eq. (9) we realize that the boundary condition (10a) should be specified as follows: In the vicinity of $u=1$
$p=s_{1}(u-1).$ (24)
For practical purposes we can sometimes continue to use this boundary condition in the previous form (10a), like we deed in Section III. The differential equation itself produces the correct condition (24). However, if we want, for example, to solve (9) numerically, the condition (24) should be used from the beginning.
Now let us come to (17). The characteristic equation for the system is
$s^{2}+cs+F^{\prime}(0)=0.$ (25)
The roots of the characteristic equations are
$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4F^{\prime}(0)}}{2}.$ (26)
We have to consider separately the case of the real roots, the case of the complex roots and the case of the double root.
Let us start from the case of the real roots ($c^{2}>4F^{\prime}(0)$). In this case we can reduce the system (17) to
$\frac{dX}{s_{3}X}=\frac{dY}{s_{4}Y},$ (27)
where
$X$ $=p+s_{4}v,$ (28a)
$Y$ $=p+s_{3}v.$ (28b)
The general integral is given by the equation
$Y=CX^{\frac{s_{4}}{s_{3}}}.$ (29)
Because $s_{3}$ and $s_{4}$ have the same sign, $X$ approaches zero with $Y$, and all the integral curves pass through the origin, which is a node. (Note that $c$ given by (14) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
For $c^{2}<4F^{\prime}(0)$ the characteristic equation has two conjugate imaginary roots $\alpha+\beta i,\alpha-\beta i$, $\beta\neq 0$) we can reduce the system (17) to the form goursat
$\frac{dX}{\alpha X-\beta Y}=\frac{dY}{\beta X+\alpha Y}$ (30)
from which we derive
$\frac{XdX+YdY}{\alpha\left(X^{2}+Y^{2}\right)}=\frac{XdY-YdX}{\beta\left(X^{2}+Y^{2}\right)}$ (31)
The general integral of the equation (17) is therefore represented by the equation
$\sqrt{X^{2}+Y^{2}}=Ce^{\frac{\alpha}{\beta}\tan^{-1}\frac{Y}{X}}.$ (32)
All these curves have the form of spirals which approach the origin as an asymptotic point. The origin is said to he a focus.
If the characteristic equation has a double root $s=c/2$, which occurs for $c^{2}=4F^{\prime}(0)$, the system (17) reduces to the form
$\frac{dX}{X}=\frac{dY}{X+Y},$ (33)
and the general integral is $Y=CX+X\ln X$. In order to construct these curves, we can express $X$ and $Y$ in terms of an auxiliary variable by putting $X=e^{\theta}$, which gives $Y=Ce^{\theta}+\theta e^{\theta}$. When $\theta$ approaches $\infty$- ∞, $X$ and $Y$, and consequently $x$ and $y$, approach zero, and the origin is again a focus. Note that if for physical reasons $u$ should be strictly nonnegative, the focus solutions (and hence the value of speed $c$ for which $c^{2}\leq 4F(0)$) are unacceptable.
Now we understand that the second boundary condition (10b) can be discarded. It is satisfied automatically. And returning to the previous Section: we did finetune the speed $c$ for the solution of (11b) to be expressed in terms of elementary functions, and not for the solution to exist.
Appendix A Generalized Burgers–Huxley equation
The results obtained in Section II for the generalized KPP equation (11a) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (34)
where $\lambda$ is a constant. (The Burgers–Huxley equation is known in various fields of applied mathematics. For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too.)
For the travelling wave solution of (34) we obtain instead of (11b) the equation
$pp_{u}+\lambda u^{n}p+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (35)
One can easily see that the solution (12) (and hence Eq. (15)) is valid in the present case also, same as the formula (14). The only difference is that Eq. (13) in this case takes the form
$P=\frac{\sqrt{\lambda^{2}+4(n+1)}-\lambda}{2(n+1)}.$ (36)
If we put $a=0$ in the formula (14), Eqs. (36) and (15) supply the solution of the generalized Huxley equation clarkson
$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u^{n+1}\left(1-u^{n}\right).$ (37)
Appendix B Perturbation series
Let us return to Eq. (9). In this Appendix we’ll present the expansion of the solution of the equation (and hence of the solution of Eq. (5)) with respect to the powers of $1/c$ (meaningful if $c\gg 1$). Such expansion has the form
$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (38)
where
$p^{(1)}$ $=F(u),$ (39a)
$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u}.$ (39b)
Let us restrict ourselves by the next to leading order approximation with respect to the parameter $1/c$. Substituting thus truncated $p(u)$ into (8) we obtain
$\frac{du}{d\xi}=\frac{1}{c}F(u)\left(1-\frac{1}{c}\frac{dF(u)}{du}\right)$ (40)
or, equivalently (with the accepted precision),
$\frac{d\xi}{c}=\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (41)
Integrating (41) we get
$\frac{\xi}{c}=\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (42)
To make our life simpler, further on consider only the particular case of $F(u)$ corresponding to the Fisher equation:
$F(u)=u(u-1).$ (43)
Performing integration in (42) we obtain
$\frac{\xi}{c}=\ln\left[\frac{1-u}{u}\right]+\frac{1}{c}\ln\left[u(1-u)\right].$ (44)
In the leading order with respect to $1/c$ the solution of (44) is
$u(\xi)=$ $\frac{1}{\exp\left(\xi/c\right)+1},$ (45)
hence with the accepted precision we can rewrite (44) as
$\frac{\xi}{c}-\frac{1}{c}\ln\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}=\ln\left[\frac{1-u}{u}\right].$ (46)
Exponentiating, solving with respect to $u$ and again expanding with respect to $1/c$ we obtain after a bit of algebra
$u(\xi)=\frac{1}{\exp\left(\xi/c\right)+1}$
$+\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$ (47)