### v5

# The exact travelling wave solutions of a KPP equation

Abstract
We obtain the exact analytical traveling wave solutions of the Kolmogorov-Petrovskii-Piskunov equation with the reaction term belonging to the class of functions, which includes that of the (generalized) Fisher equation, for the particular values of the waves speed. Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
I Introduction
Fisher fisher first introduced a nonlinear evolution equation to investigate the wave propagation of an advantageous gene in a population. His equation also describes the logistic growth–diffusion process and has the form
$u_{t}-\nu u_{xx}=ku\left(1-\frac{u}{\kappa}\right)$ (1)
where $\kappa(>0)$ is the carrying capacity of the environment. The term $f(u)=ku(1-u/\kappa)$ represents a nonlinear growth rate which is proportional to $u$ for small $u$, but decreases as $u$ increases, and vanishes when $u=\kappa$. Equation (1) arises in many physical, biological, and chemical problems involving diffusion and nonlinear growth. It is convenient to introduce the nondimensional quantities $x^{*},t^{*},u^{*}$, so after dropping the asterisks, equation (1) takes the nondimensional form
$u_{t}-u_{xx}=u\left(1-u\right).$ (2)
In the same year (1937) as Fisher, Kolmogorov, Petrovsky and Piskunov (KPP) kolmogorov introduced the more general reaction-diffusion equation, which in the nondimensional form can be presented as
$u_{t}-u_{xx}=F(u),$ (3)
where $F$ is a sufficiently smooth function with the properties that $F(0)=F(1)=0$, $F^{\prime}(0)>0$ and $F^{\prime}(1)<0$.
Equation (3) admits traveling wave solutions
$u(x,t)=u(\xi),\hskip 14.22636pt\xi=x-ct,$ (4)
where $c$ is the wave speed, and $u(\xi)$ satisfies the equation
$u_{\xi\xi}+cu_{\xi}=F(u)$ (5)
with the boundary conditions
$\lim_{\xi\to-\infty}u(\xi)$ $=1,$ (6a)
$\lim_{\xi\to+\infty}u(\xi)$ $=0.$ (6b)
Equation (5) has a simple mechanical interpretation. It describes motion of the Newtonian particle with the mass 1 ($u$ being the coordinate of the particle and $\xi$ the time) in the potential well $U(u)$ defined by the equation
$\frac{dU}{du}=-F(u).$ (7)
The motion is with friction, and the coefficient of friction is equal to $c$. The point $u=0$ is the point of stable equilibrium, and the point $u=1$ unstable.
Equation (5) is a nonlinear nonintegrable differential equation. The Cauchy problem cannot be solved by the inverse scattering transform for this equation and the problem of finding some exact solutions for this equation is an important task. Several methods of obtaining such solutions has been developped. We will mention here a few of them: the tanh-function method 4 ; 5 ; 6 , the Jacobi elliptic function expansion method 8 , the simplest equation method 11 ; 12 , the Exp-function method 14 , the $G^{\prime}/G$ expansion method 15 ; 16 ; 17 ; 18 , and the Kudryashov method 19 .
In this paper we will use the method of exclusion of the independent variable benguria ; kogan3 ; kogan4 ; kogan5 , which allows to reduce the problem of integration of the second order ODE to the problem of integration of the first order ODE, to obtain the exact solutions of Eq. (5). Thus we were able to obtain the exact solutions of the KPP equation which to the best of our knowledge were not known before. Additionally we obtain the the exact analytical traveling wave solutions of the generalized Burgers–Huxley equation.
II The ODE which doesn’t contain explicitly the independent variable
Equation (5) doesn’t contain explicitly the independent variable $\xi$. This prompts the idea to consider $u$ as the new independent variable and
$p=\frac{du}{d\xi}$ (8)
as the new dependent variable. In the new variables (5) takes the form of Abel equation of the second kind polyanin .
$pp_{u}+cp=F(u).$ (9)
The boundary conditions for (5) were (6). The boundary conditions for (9) are (however, see the next Section)
$p(1)$ $=0,$ (10a)
$p(0)$ $=0.$ (10b)
Now let us be more specific and consider
$F(u)=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (11)
where $n>0$ and $a\geq 0$ are some constants, so (5) takes the form
$u_{\xi\xi}+cu_{\xi}=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (12)
Note that for $n=1/2$ and $a=1$ we recover Fisher equation. Equation (12) with $a=1$ but $n\neq 1/2$ we may call the generalized Fisher equation kaliappan ; murray ; debnath .
Starting from (12), we obtain Eq. (9) in the form
$pp_{u}+cp=u\left(1-u^{n}\right)\left(u^{n}+a\right).$ (13)
One can easily check up that there exists the solution of (13) satisfying the boundary conditions (10)
$p(u)=p_{0}(u)\equiv Pu\left(u^{n}-1\right),$ (14)
where
$P=\frac{1}{\sqrt{n+1}},$ (15)
provided that the speed of the wave is
$c=c_{0}\equiv P+\frac{a}{P}.$ (16)
Substituting thus obtained $p(u)$ into (8) and integrating we obtain the solution of (12) as
$u(\xi)=\frac{1}{\left[\exp(nP\xi)+1\right]^{1/n}}.$ (17)
(An alternative way to obtain (17) was presented in our previous publication kogan4 .)
For the Fisher equation from (17) follows the well known result ablo
$u(\xi)=\frac{1}{\left[\exp\left(\xi/\sqrt{6}\right)+1\right]^{2}},$ (18)
and the condition (16) takes the form $c=5/\sqrt{6}$.
In Appendix A we’ll see how the results (14) - (17) can be generalized in the case when advection is taken into account in addition to diffusion. In Appendix C we’ll analyze the solution of (13) (for $n=1$) for $c$ close to $c_{0}$.
III The correct boundary conditions for Eq. (9)
In the previous Section we postulated two boundary conditions (10) for (9). However, the equation is of the first order. The first impression is that the problem is overdetermined, and the solution obtained there exists due to the finetuning of the l.h.s. of the equation. So one may think that we should finetune the speed $c$ for the solution of (13) to exist, and the solution will not exist for a general value of $c$. In fact, didn’t we finetune $c$ in the previous Section? However, the last statement immediately comes into contradiction with the results of the Appendix B, where the solution of (13) for arbitrary $c\gg 1$ is presented. It looks like we have a problem.
To solve this problem we have to return to Eq. (5) with the boundary conditions (6). Actually, the mechanical analogy allows us easily visualize the solution of the equation, but let us forget for the time being about this analogy and look at the equation purely mathematically. We immediately realize that the boundary conditions (6) - they are exceptional goursat $u_{\xi\xi}$ is not defined at $\xi=\pm\infty$. To make use of the boundary conditions consider Eq. (5) for $\xi\to\pm\infty$. For $\xi\to+\infty$ the equation can be presented as
$u_{\xi\xi}+cu_{\xi\xi}+F^{\prime}(0)u=0.$ (19)
For $\xi\to-\infty$ the equation can be presented as
$v_{\xi\xi}+cv_{\xi}-|F^{\prime}(1)|v=0,$ (20)
where $v=1-u$.
The solution of (20) is
$v=Ae^{s_{1}\xi}+Be^{s_{2}\xi},$ (21)
where $A$ and $B$ are arbitrary constants, and $s_{1,2}$ are the roots of the characteristic equation
$s^{2}-cs-|F^{\prime}(1)|=0,$ (22)
that is
$s_{1}$ $=\frac{c+\sqrt{c^{2}+4|F^{\prime}(1)|}}{2},$ (23a)
$s_{2}$ $=\frac{c-\sqrt{c^{2}+4|F^{\prime}(1)|}}{2}.$ (23b)
Because of the boundary condition (6a), $B=0$. So this boundary condition should be specified to: For $\xi\to-\infty$
$1-u\sim e^{s_{1}\xi}.$ (24)
Now let us come to (19). The characteristic equation is
$s^{2}+cs+F^{\prime}(0)=0,$ (25)
that is
$s_{3,4}=\frac{-c\pm\sqrt{c^{2}-4F^{\prime}(0)}}{2}.$ (26)
We have either the real roots (both being negative), the double (negative) root, or the complex roots (both with the negative real part). In the first case solution of (19) is
$u=Ae^{s_{3}\xi}+Be^{s_{4}\xi},$ (27)
for the case of double root the solution is
$u=e^{-c\xi/2}(A+B\xi),$ (28)
and in the complex case the solution is
$u=e^{-c\xi/2}\left[A\cos(\alpha\xi)+B\sin(\alpha\xi)\right],$ (29)
where $\alpha=[4F^{\prime}(0)-c^{2}]/2$. In any case the solution of (19) satisfies the boundary condition (6b) for any $A$ and $B$. So the boundary condition (6b) can be discarded.
Because we have lost one boundary condition the solution of (5) contains an arbitrary constant. And here our mechanical analogy helps us to understand that this (irrelevant) constant describes the translation of the solution in ”time” $\xi$.
After we understood the true nature of the boundary conditions (6) for (5), the situation with the boundary conditions (10) for Eq. (9) becomes obvious. Note that the boundary conditions (10) are also exceptional goursat $p_{u}$ is not defined either at $u=0$ or at $u=1$.
In the vicinity of $u=1$, Eq. (9) can be simplified to
$qq_{v}-cq=|F^{\prime}(1)|v,$ (30)
where again $v=1-u$, and $q\equiv dv/d\xi$. Equation (30) has two solutions
$q$ $=s_{1}v,$ (31a)
$q$ $=s_{2}v,$ (31b)
where $s_{1,2}$ are given by (23). Only the solution (31a) satisfies the condition (6a). Hence we realize that the boundary condition (10a) should be specified as follows: For $u\to 1$
$p=s_{1}(u-1).$ (32)
(Not the difference between the proportionality sign in (24) and the equality sign in (32).) One can easily check up that our solution (14) satisfies this boundary condition.
In the vicinity of $u=0$, Eq. (9) can be simplified to
$pp_{v}+cp=F^{\prime}(0)u.$ (33)
Equation (33) has two solutions
$p$ $=s_{3}u,$ (34a)
$p$ $=s_{4}u,$ (34b)
where $s_{3,4}$ are given by (26). Because both $s_{3}$ and $s_{4}$ have negative real part, both solutions will satisfy the boundary condition (6b). So we understood that the second boundary condition (10b) can be discarded – it is satisfied automatically.
So finally, in the previous Section we were solving the first order ODE with the single boundary condition. The solution should exist for any given value of the speed $c$ and shouldn’t contain an undefined constant. We did finetune the speed $c$ in the previous Section for the solution of (13) to be expressed in terms of elementary functions, and not for the solution to exist.
Our analysis of (19) allows us to understand (partially) the analytic properties of $p(u)$. If $c^{2}\geq 4F^{\prime}(0)$, from (27) and (28) we can expect that $p(u)$ would be either a single - or a double-valued function. (Further analysis is necessary to decide which option is realised.) Note that $c$ given by (16) obviously fulfills the inequality $c^{2}>4a$ for any $n>0$ and $a\geq 0$.)
Looking at (29) (which corresponds to $c^{2}<4F^{\prime}(0)$) on the other hand, we realize that $p(u)$ would be an infinitely-valued function. Note that if for physical reasons $u$ should be strictly nonnegative, such values of speed are unacceptable, because from (29) follows that $u$ would acquire negative values in the vicinity of $\xi=+\infty$. (Apart from recovering this well known property, we in this paper will say nothing about the front propagation speed selection, an issue which is both important and complicated saarloos ; ma .)
Acknowledgements.
We are grateful to G. G. Contreras, A. Gupta, M. Ma, B. Malomed, C. Ou and R. Selvaraj for the insightful discussions.
Appendix A Generalized Burgers–Huxley equation
The results obtained in Section II for the generalized KPP equation (12) can be easily generalised to the generalized Burgers–Huxley equation cherniha ; kushner
$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u\left(1-u^{n}\right)\left(u^{n}+a\right),$ (35)
where $\lambda$ is a constant. The Burgers–Huxley equation is known in various fields of applied mathematics. For example, it describes transport processes in systems when diffusion and convection are equally important and nonlinear reaction–diffusion processes too. (An interesting further generalization of the model see in gupta .)
For the travelling wave solution of (35) we obtain instead of (13) the equation
$pp_{u}+\lambda u^{n}p+cp=u\left(u^{n}-1\right)\left(u^{n}+a\right).$ (36)
One can easily see that the solution (14) (and hence Eq. (17)) is valid in the present case also, same as the formula (16). The only difference is that Eq. (15) in this case takes the form
$P=\frac{\sqrt{\lambda^{2}+4(n+1)}-\lambda}{2(n+1)}.$ (37)
If we put $a=0$ in the formula (16), Eqs. (37) and (17) supply the solution of the generalized Huxley equation clarkson
$u_{t}-\lambda u^{n}u_{x}-u_{xx}=u^{n+1}\left(1-u^{n}\right).$ (38)
Appendix B Perturbation theory I
Let us return to Eq. (9). In this Appendix we’ll present the expansion of the solution of the equation (and hence of the solution of Eq. (5)) with respect to the powers of $1/c$ (meaningful if $c\gg 1$). Such expansion has the form
$p(u)=\sum_{k=1}\frac{1}{c^{k}}p^{(k)}(u),$ (39)
where
$p^{(1)}$ $=F(u),$ (40a)
$p^{(k)}$ $=-\sum_{m=1}^{k-1}p^{(m)}p^{(k-m)}_{u},\hskip 14.22636ptk=2,3,\dots.$ (40b)
Let us restrict ourselves by the next to leading order approximation with respect to the parameter $1/c$. Substituting thus truncated $p(u)$ into (8) we obtain
$\frac{du}{d\xi}=\frac{1}{c}F(u)\left(1-\frac{1}{c}\frac{dF(u)}{du}\right)$ (41)
or, equivalently (with the accepted precision),
$\frac{d\xi}{c}=\frac{du}{F(u)}+\frac{1}{c}\frac{dF(u)}{F(u)}.$ (42)
Integrating (42) we get
$\frac{\xi}{c}=\int\frac{du}{F(u)}+\frac{1}{c}\ln|F(u)|.$ (43)
To make our life simpler, further on consider only the particular case of $F(u)$ corresponding to the Fisher equation:
$F(u)=u(u-1).$ (44)
Performing integration in (43) we obtain
$\frac{\xi}{c}=\ln\left[\frac{1-u}{u}\right]+\frac{1}{c}\ln\left[u(1-u)\right].$ (45)
In the leading order with respect to $1/c$ the solution of (45) is
$u(\xi)=$ $\frac{1}{\exp\left(\xi/c\right)+1},$ (46)
hence with the accepted precision we can rewrite (45) as
$\frac{\xi}{c}-\frac{1}{c}\ln\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}=\ln\left[\frac{1-u}{u}\right].$ (47)
Exponentiating, solving with respect to $u$ and again expanding with respect to $1/c$ we obtain after a bit of algebra
$u(\xi)=\frac{1}{\exp\left(\xi/c\right)+1}$
$+\frac{1}{c}\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\ln\left\{\frac{\exp\left(\xi/c\right)}{\left[\exp\left(\xi/c\right)+1\right]^{2}}\right\}.$ (48)
Appendix C Perturbation theory II
In Section II we have solved Eq. (13) for $c=c_{0}$. Now let us study the solution for $c$ close to $c_{0}$, considering $\delta c\equiv c-c_{0}$ as the small parameter and using the perturbation series. To make life simpler $n=1$, we’ll consider only the case $n=1$. Presenting the solution as
$p(u)=p_{0}(u)+\frac{\delta c}{\sqrt{2}}\delta p(u),$ (49)
for $\delta p(u)$ in the lowest with respect to $\delta c$ order we obtain equation
$u(u-1)\delta p_{u}+\left(2u-1+c_{0}\right)\delta p+u(u-1)=0.$ (50)
Integrating (50) we obtain
[…]
where
[…]
and $C$ is an arbitrary constant. The lower limits of the integrals in (51) and (52) can be taken arbitrary. Calculating the integral in (52) we get
$\delta p(u)=\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\left(C-\int\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}\right).$ (53)
It will be convenient to choose the lower limit in the in integral in (53) to be equal to zero. After that, for $\delta p$ to remain finite when $u\to 0,1$ we should choose $C=0$, thus obtaining
$\delta p(u)=-\frac{u^{c_{0}-1}}{\left(1-u\right)^{c_{0}+1}}\int_{0}^{u}\frac{\left(1-u\right)^{c_{0}+1}du}{u^{c_{0}-1}}.$ (54)
A pleasant surprise is that $\delta p(u)$, as given by (54), goes to zero when $u\to 0,1$. (We couldn’t be sure in that just looking at Eq. (50), because the coefficient before $\delta p_{u}$ in the equation goes to zero in the above mentioned limits, hence one can doubt the validity of the linear approximation with respect to $\delta c$ in the vicinity of $u=0$ and $u=1$. )